# Automatic Stage Lighting Control: Is it a Rule-Driven Process or Generative Task?

### v1

@@ Abstract.
Stage lighting plays an essential role in live music performances, influencing the engaging experience of both musicians and audiences. Given the high costs associated with hiring or training professional lighting engineers, Automatic Stage Lighting Control (ASLC) has gained increasing attention. However, most existing approaches only classify music into limited categories and map them to predefined light patterns, resulting in formulaic and monotonous outcomes that lack rationality. To address this issue, this paper presents an end-to-end solution that directly learns from experienced lighting engineers – Skip-BART. To the best of our knowledge, this is the first work to conceptualize ASLC as a generative task rather than merely a classification problem. Our method modifies the BART model to take audio music as input and produce light hue and value (intensity) as output, incorporating a novel skip connection mechanism to enhance the relationship between music and light within the frame grid. We validate our method through both quantitative analysis and an human evaluation, demonstrating that Skip-BART outperforms conventional rule-based methods across all evaluation metrics and shows only a limited gap compared to real lighting engineers. Specifically, our method yields a p-value of 0.72 in a statistical comparison based on human evaluations with human lighting engineers, suggesting that the proposed approach closely matches human lighting engineering performance. To support further research, we have made our self-collected dataset, code, and trained model parameters available at https://github.com/RS2002/Skip-BART.
Automatic Stage Lighting Control (ASLC), Music Information Retrieval (MIR), Multi-Modal, Generation Model
1. Introduction
@@
Recently, multi-modal machine learning has significantly advanced various fields in Music Information Retrieval (MIR), including video-to-music generation (Tian et al., 2024), music video (MV) generation (Chen et al., 2025), and lyrics generation (Ma et al., 2024). Among these techniques, Automatic Stage Lighting Control (ASLC)  [^1] To avoid confusion, in this paper, ”lighting control” and ”lighting generation” are used interchangeably. plays a crucial role in live performances by shaping the atmosphere and influencing the emotions of both musicians and audiences, while offering an affordable cost. Most existing ALSC methods are based on predefined rules that classify music into finite categories with deterministic patterns. The style-based (Stanescu et al., 2018) and emotion-based (Hsiao et al., 2017) classification methods are two mainstream approaches. However, these methods have several limitations: First, due to limited datasets and labeling difficulty, the accuracy of classification models is often insufficient, and the defined categories are typically coarse-grained. Besides the biases caused by misclassification, this coarse-grained labeling directly impacts model performance. For example, the atmospheres of black metal and folk black metal are entirely different, yet they may be grouped with thrash metal under the metal category, or even combined with punk in the broader rock category. Second, as lighting is viewed as an amalgam of technology and art (Gillette and McNamara, 2019), it is debatable whether we can simply use certain patterns to define it. Additionally, McDonald et al. (McDonald et al., 2022) have demonstrated that limited relationship between hue of light and emotion, raising questions about the appropriateness of mapping emotions solely to the light.
@@
To address this research gap, we propose to treat ALSC as an art content generation task rather than a mechanized mapping process. Inspired by the success of generative models in the field of MIR, we develop an end-to-end deep learning methodology to train a network directly on ground truth data from real lighting engineers. To begin with, we introduce an automatic label generation method from video data and create a self-collected dataset to address the absence of training dataset in the field. Next, we introduce a transfer learning and pre-training method to enable the model to learn effectively from limited data. We evaluate our methodology through both quantitative analysis and human evaluation. The result shows that our method outperforms conventional rule-based approaches across all metrics and achieves comparable performance to that of human lighting engineers.
In summary, the main contributions of this paper are as follows:
  • Methodology: We develop Skip-Bidirectional Auto-Regressive Transformers (Skip-BART), a deep learning method for ALSC. In terms of model structure, we modify the embedding layers of the original BART to support both audio and light inputs. Additionally, we introduce a novel skip-connection mechanism to enhance the relationship between music and light within the frame-grained context. For the training method, we introduce a transfer learning and pre-training approach to enable the model to learn sufficiently from limited data. A restricted stochastic temperature-controlled sampling method is also proposed to maintain both diversity and stability in the generation results. To the best of our knowledge, this is the first end-to-end method that directly learns from human lighting engineers for ALSC.
  • Dataset: Given the lack of related datasets for ALSC, we introduce a label generation method from video data. Following this method, we have collected and released a stage lighting dataset titled Rock, Punk, Metal, and Core - Livehouse Lighting (RPMC-L2), which is the first dataset for stage lighting.
@@
  • Experiment: Through quantitative analysis and human evaluation, we demonstrate that our method outperforms conventional rule-based approaches across all metrics. Specifically, our method yields a significant difference ($p<0.001$) compared to conventional rule-based methods, while showing no significant difference compared to human lighting engineers ($p=0.72$) in the human evaluation. This suggests that our proposed method closely matches human lighting engineering performance.
2. Related Work
2.1. Automatic Stage Lighting Control
@@
Although ALSC is not a widely explored topic in MIR, it has received some attention in previous works. Most prior methods can be summarized as a paradigm that first classifies music into predefined categories, such as emotion and style, and then defines a corresponding light pattern for each category. For instance, works (Stanescu et al., 2018; Lei et al., 2021; Kanno and Fukuhara, 2022; Liao et al., 2023) utilize style and theme for classification based on audio or lyrics, respectively, while Mabpa et al. (Mabpa et al., 2021) classify music using chords. However, these methods exhibit significant limitations. Firstly, there is insufficient theoretical support for mapping these categories to light. Additionally, coarse-grained classification severely impacts model performance. For example, Mabpa et al. (Mabpa et al., 2021) classify chords into only 24 major and minor chords, neglecting more nuanced chords, such as augmented, seventh, and suspended chords, which possess distinct colors and functions. Furthermore, many pieces of music lack lyrics or contain lyrics that are difficult to recognize (such as screaming in heavy music), rendering the method of Liao et al. (Liao et al., 2023) ineffective.
@@
Additionally, a series of works focus on emotion classification (Bonde et al., 2018; Hsiao et al., 2017; Moon et al., 2015), employing various machine learning methods, including Support Vector Machines (SVM), Support Vector Regression (SVR), and Multi-Layer Perceptrons (MLP). Vryzas et al. (Vryzas et al., 2017) extend this to the field of drama. However, the relationship between music or performance emotion and light remains controversial, with different studies presenting varying perspectives (Kanno and Fukuhara, 2022; McDonald et al., 2022). We argue that simply mapping emotion to light is not a viable approach for two reasons: (i) Although Kanno et al. (Kanno and Fukuhara, 2022) demonstrate that certain light colors are frequently associated with specific emotions by professional lighting engineers, this does not imply that these colors must be used when conveying those emotions. For example, in the dataset associated with this paper, we observe that the green hue appears in groove metal associated with anger emotion, contradicting the color theory suggesting red should be used. (ii) Beyond the granularity issue mentioned earlier, the performance of all rule-based methods depends on the accuracy of the classifier. In some works (Zhao, 2025), the emotion classification accuracies of various methods fall below 80%, raising doubts about the practicality of these methods at their current stage.
@@
Moreover, Tyroll et al. (Tyroll et al., 2020) present a different workflow, where a feature extractor is first trained via an autoencoder, and then the feature embeddings are directly mapped to light. However, the feature extraction capacity of autoencoders has been shown to be limited (Van Den Oord et al., 2017), and the mapping from feature embeddings to light is also ambiguous. In contrast, we propose the first end-to-end method for stage lighting control, framing the task as a generation problem rather than a classification task. Compared to previous hierarchical mapping methods, our approach directly trains the model on real-world data from lighting engineers, resulting in improved performance.
2.2. Transformer in Music Information Retrieval
Music exhibits a high degree of similarity to natural language, particularly in terms of contextual relationships. Numerous studies have demonstrated the effectiveness of Transformers (Vaswani, 2017) in the field of MIR. For instance, encoder-only Transformers with bidirectional attention have shown promising results in music understanding tasks (Zhu et al., 2021), while decoder-only Transformers with unidirectional attention excel in music generation tasks (Du et al., 2023). The encoder-decoder Transformer combines the advantages of both architectures, yielding promising outcomes in more generalized tasks (Yang et al., 2023). Notably, the encoder-decoder structure allows for the integration of different modalities within the two streams, making it particularly suitable for multi-modal applications. For example, Zhang et al. (Zhang et al., 2022) combines music and text information in the encoder and employs the decoder to generate interpretations of lyrics. Kang et al. (Kang et al., 2024) utilizes video as the input to the encoder, while incorporating music information such as chords and key signatures in the decoder to generate appropriate music for a video.
Furthermore, unsupervised pre-training has not only led to significant advancements in Large Language Models (LLMs) but has also proven beneficial in enabling models to learn more deeply from limited music datasets. For example, Zhao (Zhao, 2025) proposed an adversarial Mask Language Model (MLM) approach to help the model understand the underlying structure of effective music data. Liang et al. (Liang et al., 2024) and Zeng et al. (Zeng et al., 2021) developed various multi-level object masking techniques to prevent information leakage, which have resulted in substantial improvements in downstream tasks. In this paper, we present an encoder-decoder Transformer that takes audio music as input for the encoder and uses the decoder to generate the corresponding lighting sequence in an auto-regressive manner. Moreover, we adapt the MLM pre-training task to enable the model to learn the structure of music independently.
3. Methodology
3.1. Overview
@@
Figure 1. Network Architecture: In the figure, ‘ice’ represents the frozen parameters, and ‘fire’ denotes the trainable ones.
In this section, we introduce an end-to-end model, Skip-BART, to generate single principal stage lighting based on music. Similar to previous works, we focus solely on offline primary single lighting generation. We first analyze the entire audio track, generating only a single hue and value for each frame to represent the overall characteristics of the stage light. Since we only consider the single light condition, we treat it as a pure color and fix the saturation in HSV color space at 100%.
@@
In the following content, we first introduce our model structure, as shown in Fig. 1. We modify the embedding layers to support our task’s input, utilize pre-trained models from other tasks for model enhancement based on transfer learning, and incorporate a skip-connection mechanism to better capture the aligned relationship between music and light. Subsequently, we outline the workflow of our method, which includes MLM-based pre-training, end-to-end fine-tuning, and inference using a restricted stochastic temperature-controlled sampling method.
3.2. Data Process
Before delving into the details of our model, we first describe how we obtain the audio and image sequences from raw video to support our end-to-end training. First, we divide the video into audio sequences $s_{a}$ and image sequences $s_{p}$ at the same sampling rate, ensuring that both modalities have the same number of frames $T$.
For audio sequences $s_{a}$, we use feature extractor to obtain the music sequence $x=f_{\text{audio}}(s_{a})$. Here, we select the pre-trained Open L3 (Cramer et al., 2019) as the feature extractor $f_{\text{audio}}$, which has proven effective in numerous audio-related studies (Wilkinghoff, 2021; Koh and Dubnov, 2021).
To extract the primary light from each image, we first convert each frame into the HSV color space. After flattening, each frame yields a set of $(h,s,v)$ tuples corresponding to its pixels. As described previously, we fix the saturation component $s$, and therefore omit it in subsequent description. We focus on the hue and value components and compute the distributions $D_{h,t}$ and $D_{v,t}$ for frame $t$ as follows:
(1) $D_{h,t}$ $=\text{Histogram}\left(\left\{h_{t}^{k}\mid v_{t}^{k}\geq v^{\prime}\right\}\right),$
$D_{v,t}$ $=\text{Histogram}\left(\left\{v_{t}^{k}\mid v_{t}^{k}\geq v^{\prime}\right\}\right),$
where $h_{t}^{k}$ and $v_{t}^{k}$ denote the hue and value components, respectively, of the $k$-th pixel in the $t$-th frame, and $v^{\prime}$ is a threshold used to avoid hue distortion caused by extreme light intensity.
(2) $y_{t}=[\overline{h_{t}},\overline{v_{t}}]\in\mathbb{R}^{2}_{+},\quad\text{where}\quad\begin{cases}\overline{h_{t}}=\operatorname{Mode}(D_{h,t})\\ \overline{v_{t}}=\dfrac{\sum_{k=1}^{T}k\cdot D_{v,t}(k)}{\sum_{k=1}^{T}D_{v,t}(k)}\end{cases}$
We use the mode to compute the tokenized hue, since hue is a circular variable.
This results in a music sequence $x=\{x_{1},x_{2},\ldots,x_{T}\}$ and the corresponding light sequence $y=\{y_{1},y_{2},\ldots,y_{T}\}$. We consider the music sequence $x$ as input and the corresponding light sequence $y$ as output.
Our data processing method is compared with the direct extraction approach in Fig. 2. Our method enables more accurate color extraction under extreme intensity lighting conditions.
@@ 3.2. Data Process
Figure 2. Each subfigure shows the input scene on the top, the result of the direct extraction method in the bottom-left, and our extraction method in the bottom-right. (a)-(b) Both methods accurately extract the dominant hue. (c)-(d) Our method extracts colors that are closer to the original appearance.
3.3. Network Architecture
The model structure is illustrated in Fig. 1. In our task, the model takes an entire piece of music as input and generates the corresponding lighting sequence. We believe that BART (Lewis, 2019) is well-suited for this task, as the bi-directional encoder effectively extracts contextual information simultaneously, while the unidirectional decoder generates outputs in an auto-regressive manner, thereby maintaining correct contextual correlations. To adapt BART for our task, we first modify its embedding layers to support the music and light input. To align the dimensions between Open L3 and BART, we use a MLP to further extract features from the embeddings produced by Open L3. The hue and value obtained from Section 3.2 are then embedded using respective embedding layers and subsequently combined, allowing them to be input into the decoder of the backbone. As noted in (Zhao et al., 2024b), the tokenization process may lead to some information loss. However, directly processing hue data with an MLP presents challenges due to the cyclical nature of hue values: the minimum and maximum hue values can be quite similar, complicating the learning process. In contrast, the embedding layer effectively captures the relationships between different tokens, making it more suitable for handling cyclic data like hue.
@@ 3.3. Network Architecture
Additionally, transfer learning has demonstrated the benefits of using trained parameters from other domains (Pan, 2014), and also shown the efficiency in signal processing fields (Zhao et al., 2025, 2024a). Given the limited datasets in MIR, we aim to introduce more external knowledge to enhance model performance. Consequently, we adopt the backbone of PianoBART (Liang et al., 2024), a large-scale pre-trained model for symbolic music understanding and generation, in our Skip-BART framework. Additionally, to embed as much knowledge as possible, we utilize the Drop And Rescale (DARE) (Yu et al., 2024) method to combine the fine-tuned parameters of PianoBART across all downstream tasks, including priming, melody extraction, velocity prediction, composer classification, and emotion classification. The combined model parameters are expressed as:
(3) $\theta$ $=\theta_{pre}+\lambda\sum_{i}(\theta^{i}_{DARE}-\theta_{pre})\ ,$
@@
where $\theta$ represents the combined model parameters, $\theta_{pre}$ denotes the pre-trained model parameters, $\theta^{i}$ represents the fine-tuned model parameters for the $i^{th}$ downstream task, $d$ is the dimension of the parameters, $p$ is the drop rate, $\lambda$ is the scaling term, and $\odot$ represents element-wise multiplication. We then employ LoRA (Hu et al., 2021) for efficient fine-tuning in our tasks. Due to page limitations, further details can be found in (Yu et al., 2024). Finally, for both pre-training and fine-tuning, we employ different heads for various tasks, which will be detailed in the next section.
3.4. Workflow
Figure 3. The workflow of Skip-BART consists of three phases: pre-training, fine-tuning, and inference.
3.4.1. Pre-training

### v2

@@ Abstract
Stage lighting is a vital component in live music performances, shaping an engaging experience for both musicians and audiences. In recent years, Automatic Stage Lighting Control (ASLC) has attracted growing interest due to the high costs of hiring or training professional lighting engineers. However, most existing ASLC solutions only classify music into limited categories and map them to predefined light patterns, resulting in formulaic and monotonous outcomes that lack rationality. To address this gap, this paper presents Skip-BART, an end-to-end model that directly learns from experienced lighting engineers and predict vivid, human-like stage lighting. To the best of our knowledge, this is the first work to conceptualize ASLC as a generative task rather than merely a classification problem. Our method adapts the BART model to take audio music as input and produce light hue and value (intensity) as output, incorporating a novel skip connection mechanism to enhance the relationship between music and light within the frame grid. To address the lack of available datasets, we create the first stage lighting dataset, along with several pre-training and transfer learning techniques to improve model training with limited data. We validate our method through both quantitative analysis and an human evaluation, demonstrating that Skip-BART outperforms conventional rule-based methods across all evaluation metrics and shows only a limited gap compared to real lighting engineers. The code and dataset of this paper are provided at https://github.com/RS2002/Skip-BART.
1 Introduction
Recently, multi-modal machine learning has significantly advanced various fields in Music Information Retrieval (MIR), including video-to-music generation (tian2024vidmuse), music video (MV) generation (chen2025mv), and lyrics generation (ma2024keyric). Among these techniques, Automatic Stage Lighting Control (ASLC)  [^1] To avoid confusion, in this paper, “lighting control” and “lighting generation” are used interchangeably. shows potential in live performances by shaping the atmosphere and influencing the emotions of both musicians and audiences, while offering an affordable cost.
@@ 1 Introduction
Currently, most existing ASLC methods are based on predefined rules that classify music into finite categories with deterministic patterns. The style-based (stanescu2018automatic) and emotion-based (hsiao2017methodology) classification methods are two mainstream approaches. However, these methods have several limitations: First, due to limited datasets and labeling difficulty, the accuracy of classification models is often insufficient, and the defined categories are typically coarse-grained. Besides the biases caused by misclassification, this coarse-grained labeling directly impacts model performance. For example, the atmosphere black metal and folk black metal are entirely different, yet they may be grouped with thrash metal under the metal category, or even combined with punk in the broader rock category. Second, as lighting is viewed as an amalgam of technology and art (gillette2019designing), it is debatable whether we can simply use certain patterns to define it. Additionally, mcdonald2022illuminating have demonstrated that limited relationship between hue of light and emotion, raising questions about the appropriateness of mapping emotions solely to the light.
@@
To address this research gap, we propose to treat ASLC as an art content generation task rather than merely a mechanized mapping process. Inspired by the success of generative models in MIR, we develop an end-to-end deep learning methodology that trains a network directly on ground truth data from professional lighting engineers. Initially, we introduce an automatic label generation method from video data and create a self-collected dataset titled *Rock, Punk, Metal, and Core - Livehouse Lighting (RPMC-L2) to address the scarcity of training datasets in this field. Subsequently, we present a carefully designed neural network called Skip-Bidirectional Auto-Regressive Transformers (Skip-BART) for stage lighting generation from music. We also incorporate transfer learning and a pre-training method to enable the model to learn effectively from limited data. We evaluate our methodology through both quantitative analysis and human evaluation. The results indicate that our method outperforms conventional rule-based approaches across all metrics and achieves performance comparable to that of human lighting engineers. In summary, the main contributions of this paper are as follows:
  • To the best of our knowledge, we are the first to frame ASLC as a generation task rather than a rule-driven mapping process, based on the intuition that ASLC is essentially an art creation process for humans. To support our claim, we study the offline primary ASLC task, which is modeled by taking a music sequence as input and generating a single light hue and value as output. We further evaluate this model on a real-world dataset, demonstrating the superiority of generative methods over previous rule-based approaches. Our novel perspective on generation provides a strong foundation for future research in ASLC.
  • We develop Skip-BART, a novel end-to-end deep learning framework to address the aforementioned generative task for ASLC. Building on BART, our approach incorporates adapted embedding and head layers to support both music and light modalities, as well as a novel skip-connection module to enhance the relationship between music and light within a fine-grained context. Additionally, we introduce pre-training and transfer learning mechanisms to improve model performance under limited training data. During inference, we implement a Restricted Stochastic Temperature-Controlled (RSTC) sampling method to ensure both diversity and stability in the generated results.
@@
  • To validate our method, we present the first stage lighting dataset, RPMC-$L^{2}$, and train Skip-BART on it. Through quantitative analysis and human evaluation, we demonstrate that our method outperforms conventional rule-based approaches across all metrics. Specifically, our method yields a significant difference ($p<0.001$) compared to conventional rule-based methods, while showing no significant difference compared to human lighting engineers ($p=0.72$) in the human evaluation. This suggests that our proposed method closely matches human lighting engineering performance. Furthermore, our method achieves promising results in cross-domain scenarios with other music styles, significantly outperforming previous rule-based methods.
2 Related Work
@@
Although ASLC is not a widely explored topic in MIR, it has received some attention in previous works. Most prior methods can be summarized as a paradigm that first classifies music into predefined categories, such as emotion and style, and then defines a corresponding light pattern for each category. For instance, works (stanescu2018automatic; lei2021music; kanno2022automatic; liao2023automatic) utilize style and theme for classification based on audio or lyrics, respectively, while mabpa2021automatic classify music using chords. However, these methods exhibit significant limitations. Firstly, there is insufficient theoretical support for mapping these categories to light. Additionally, coarse-grained classification severely impacts model performance. For example, mabpa2021automatic classify chords into only 24 major and minor chords, neglecting more nuanced chords, such as augmented, seventh, and suspended chords, which possess distinct colors and functions. Furthermore, many pieces of music lack lyrics or contain lyrics that are difficult to recognize (such as screaming in heavy music), rendering the method of liao2023automatic ineffective.
@@
Additionally, a series of works focus on emotion classification (bonde2018auditory; hsiao2017methodology; moon2015mood), employing various machine learning methods, including Support Vector Machines (SVM), Support Vector Regression (SVR), and Multi-Layer Perceptrons (MLP). vryzas2017augmenting extend this to the field of drama. However, the relationship between music or performance emotion and light remains controversial, with different studies presenting varying perspectives (kanno2022automatic; mcdonald2022illuminating). We argue that simply mapping emotion to light is not a viable approach for two reasons: (i) Although kanno2022automatic demonstrate that certain light colors are frequently associated with specific emotions by professional lighting engineers, this does not imply that these colors must be used when conveying those emotions. For example, in the dataset associated with this paper, we observe that the green hue appears in groove metal associated with anger emotion, contradicting the color theory suggesting red should be used. (ii) Beyond the granularity issue mentioned earlier, the performance of all rule-based methods depends on the accuracy of the classifier. In some works (zhao2024adversarial), the emotion classification accuracies of various methods fall below 80%, raising doubts about the practicality of these methods at their current stage.
@@
Moreover, tyroll2020avai present a different workflow, where a feature extractor is first trained via an autoencoder, and then the feature embeddings are directly mapped to light. However, the feature extraction capacity of autoencoders has been shown to be limited (van2017neural), and the mapping from feature embeddings to light is also ambiguous. In contrast, we propose the first end-to-end method for stage lighting control, framing the task as a generation problem rather than a classification task. Compared to previous hierarchical mapping methods, our approach directly trains the model on real-world data from lighting engineers, resulting in improved performance.
3 Methodology
3.1 Overview
@@
Figure 1: Network Architecture: In the figure, ‘ice’ represents the frozen parameters, and ‘fire’ denotes the trainable ones.
@@
In this section, we introduce an end-to-end model, Skip-BART, designed to generate single principal stage lighting based on music. Similar to previous works, we focus exclusively on offline primary single lighting generation. In the following content, we describe our model structure, as depicted in Fig. 1. We modify the embedding layers to accommodate our task’s input, utilize pre-trained models from other tasks for enhancement through transfer learning, and incorporate a skip-connection mechanism to better capture the aligned relationship between music and light. Subsequently, we outline the workflow of our method, which includes MLM-based pre-training, end-to-end fine-tuning, and inference using a RSTC sampling method.
Before delving into the details of our methodology, we first provide a brief introduction to the data processing, with additional details left in Appendix C.3. To facilitate the analysis of lighting conditions, the lighting information is preprocessed in the HSV color space. The HSV color space separates the brightness (Value) information and the color (Hue) information, which makes the control and generation of lighting more intuitive, independent, and stable. Specifically, we set the Saturation (S) channel to a constant value [^2] To avoid ambiguity, we use Value or V to refer to the V component in HSV, and value to represent its numerical value here. of 255 (100%) to simulate the lighting effect and counteract environmental influences, ensuring that after factors such as air scattering, fog, or surface reflections reduce color purity, the light maintains vividness and minimizes distortions like fading or whitening. Compared to the light extraction approach that only considers directly extracting HSV features, our data processing method enables more accurate color extraction under extreme intensity lighting conditions, as shown in Fig. 2.
Then, for each music sequence $x=\{x_{1},x_{2},\ldots,x_{T}\}$, we extract the corresponding light sequence $y=\{y_{1},y_{2},\ldots,y_{T}\}$, where $T$ represents the number of frames. For each frame $t$, we define $y_{t}=[\overline{h}_{t},\overline{v}_{t}]\in\mathbb{Z}^{2}_{*}$, where $\overline{h}_{t}$ and $\overline{v}_{t}$ denote the hue and value extracted from the HSV space.
@@
Figure 2: Each subfigure shows the input scene on the top, the result of the direct HSV extraction method in the bottom-left, and our extraction method in the bottom-right. (a)-(b) Both methods accurately extract the dominant hue. (c)-(d) Our method extracts colors closer to the original appearance.
3.2 Network Architecture
The model structure is illustrated in Fig. 1. In our task, the model takes an entire piece of music as input and generates the corresponding lighting sequence. We believe that BART (lewis2019bart) is well-suited for this task, as the bi-directional encoder effectively extracts contextual information simultaneously, while the unidirectional decoder generates outputs in an auto-regressive manner, thereby maintaining correct contextual correlations. To adapt BART for our task, we first modify its embedding layers to support the music and light input. To align the dimensions between Open L3 and BART, we use a MLP to further extract features from the embeddings produced by Open L3. The hue and value obtained from Appendix C.3 are then embedded using respective embedding layers and subsequently combined, allowing them to be input into the decoder of the backbone. As noted in (zhao2024mining), the tokenization process may lead to some information loss. However, directly processing hue data with an MLP presents challenges due to the cyclical nature of hue values: the minimum and maximum hue values can be quite similar, complicating the learning process. In contrast, the embedding layer effectively captures the relationships between different tokens, making it more suitable for handling cyclic data like hue.
@@ 3.2 Network Architecture
Additionally, transfer learning has demonstrated the benefits of using trained parameters from other domains (pan2014transfer). Given the limited datasets in MIR, we aim to introduce more external knowledge to enhance model performance. Consequently, we adopt the backbone of PianoBART (liang2024pianobart), a large-scale pre-trained model for symbolic music understanding and generation, in our Skip-BART framework. Additionally, to embed as much knowledge as possible, we utilize the Drop And Rescale (DARE) (yu2024language) method to combine the fine-tuned parameters of PianoBART across all downstream tasks, including priming, melody extraction, velocity prediction, composer classification, and emotion classification. The combined model parameters are expressed as:
$\theta$ $=\theta_{pre}+\lambda\sum_{i}(\theta^{i}_{DARE}-\theta_{pre})\ ,$ (1)
@@
where $\theta$ represents the combined model parameters, $\theta_{pre}$ denotes the pre-trained model parameters, $\theta^{i}$ represents the fine-tuned model parameters for the $i^{th}$ downstream task, $d$ is the dimension of the parameters, $p$ is the drop rate, $\lambda$ is the scaling term, and $\odot$ represents element-wise multiplication. We then employ LoRA (hu2022lora) for efficient tuning in our tasks. Due to page limitations, further details can be found in (yu2024language). Finally, for both pre-training and fine-tuning, we employ different heads for various tasks, which will be detailed in the next section.
In summary, our proposed network is based on the BART backbone, utilizing OpenL3 to extract music audio features and a tokenization mechanism with an embedding layer to extract light features, effectively addressing the cyclic problem of hue. A novel skip-connection mechanism is introduced to enhance the relationship between music and light within the frame grid. Additionally, the backbone directly incorporates pre-trained parameters from PianoBART, providing a solid starting point for subsequent training, while using LoRA for efficient tuning.
3.3 Workflow
Figure 3: Workflow of Skip-BART
3.3.1 MLM-based Pre-training